OpenStudy (anonymous):

Solve 3 log 2x=4 answer in decimal form

OpenStudy (zzr0ck3r):

\(3\log_2(x)=4\iff\log_2(x) = \frac{4}{3}\iff 2^{\frac{4}{3}}=x\)

OpenStudy (anonymous):

no the log base is ten not two

OpenStudy (zzr0ck3r):

log(2x)=4/3 implies 10^(4/3)=2x

OpenStudy (cggurumanjunath):